OpenStudy (anonymous):

True or False? Because the length of a rectangle is an irrational number, the perimeter and area would always be irrational. Please explain and use examples, Thank you.

OpenStudy (ikram002p):

for a circle or rectangle or triangle ??

OpenStudy (anonymous):

Rectangle, sorry.

OpenStudy (anonymous):

P = 2L + 2W A = W x H

OpenStudy (anonymous):

Can you please give examples too?

OpenStudy (ikram002p):

I will answer for rectangle first case for area : you want to prove that if \[n \neq \frac{ a }{ b } \rightarrow n^2\neq \frac{ A }{B }\] using the contrapositive assume \[n^2=\frac{ A }{ B } \rightarrow n=\frac{ a }{ b }\] when \[n^2=\frac{ A }{ b } \] then there exist a ,b such that \[n^2=\frac{ a^2 }{ b^2 }\] thats mean \[n=\frac{ a }{ b }\] which means that is a true statment

OpenStudy (anonymous):

Oh, can you give an example?

OpenStudy (ikram002p):

noting that n is length of a rectangle

OpenStudy (anonymous):

Nevermind that.

OpenStudy (anonymous):

Can you help me solve a different question?

OpenStudy (ikram002p):

ok but did you got the ansewr or there is any thing you want me to explain ?

OpenStudy (anonymous):

No, that's okay.

OpenStudy (anonymous):

Can you help me solve this question?

OpenStudy (ikram002p):

just type you question and I will see

OpenStudy (ikram002p):

Anna is the coorect answer I will explain it to you

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

So you have to put the x into the exponent?

OpenStudy (ikram002p):

I put a circle on the wrong step

OpenStudy (anonymous):

Are those the ONLY wrong steps?

OpenStudy (anonymous):

Thank you so much.